# Isocurvature Induced Gravitational Waves at Pulsar Timing Arrays

### v1

@@ Acknowledgement
We thank Lang Liu and Soubhik Kumar for useful discussions on this work and feedback on a draft of this manuscript. This work was supported in part by the National Key R&D Program of China (2021YFC2203100), by the National Natural Science Foundation of China (12433002, 12261131497), by CAS young interdisciplinary innovation team (JCTD-2022-20), by 111 Project (B23042), by CSC Innovation Talent Funds, by USTC Fellowship for International Cooperation, and by USTC Research Funds of the Double First-Class Initiative. PD is supported by the National Natural Science Foundation of China (Grants No. T2388102). JZ is supported by the National Natural Science Foundation of China (125B1023).
@@ Appendix A Evolution Functions for Primordial Perturbations
In this appendix, we show the set of Einstein equations and Boltzmann equations that govern the evolution of primordial perturbations. In the Newtonian gauge with the notation defined in section 2, the Einstein equations can be written as [Kodama:1986fg, Kodama:1986ud, Ma:1995ey]
@@
To obtain the Boltzmann equations for density perturbations, we consider baryons are strongly coupled with the SM radiation bath (denoted as $\gamma_{*}$) and use the tight-coupling approximation: setting $\theta_{b}=\theta_{\gamma_{*}}$ and neglecting the anisotropic stress $\sigma_{\gamma_{*}}=0$ [Ma:1995ey, Bucher:1999re]. With this approximation, the evolution functions of $\theta_{b}$ and $\theta_{\gamma_{*}}$ can be combined into one equation. In the radiation dominated era ($w=1/3$), the Boltzmann equations for the density perturbations of different species read
@@
$\theta_{c}^{\prime}$ $=$ $-\mathcal{H}\theta_{\rm c}+k\Phi~,$ (A.9)
@@
For DRDI with free-streaming dark radiation, we need to add an additional DR species which contains net anisotropic stress. The complete Boltzmann hierarchy for DR involves an arbitrarily large $\ell$ moment (typically denoted as $F_{\ell}$):
@@
$F_{\rm dr,\ell}^{\prime}$ $=$ $\frac{k}{2\ell+1}\left[\ell F_{\rm dr,\ell-1}-(\ell+1)F_{\rm dr,\ell+1}\right]~~~~(\ell\geq 2),$ (A.12)
where $\delta=F_{0}$, $\theta=(3/4)kF_{1}$, $\sigma=(1/2)F_{2}$. The effects of high $\ell$ moments on observables are suppressed by powers of $k\tau$ in the superhorizon limit. The Boltzmann hierarchy can therefore be simplified by truncating it at a certain $\ell_{\rm max}$ depending on the desired precision. In this study, we choose $\ell_{\rm max}=8$ and obtain closed differential equations using $F_{\ell_{\rm max}+1}\approx(2\ell_{\rm max}+1)F_{\ell_{\rm max}}/(k\tau)-F_{\ell_{\rm max}-1}$ [Ma:1995ey] and initial conditions $F_{\ell\geq 3}=0$.
@@ Appendix B Isocurvature Initial Conditions After Neutrino Decoupling
Here we show isocurvature initial conditions after neutrino decoupling, where neutrinos are treated as free-streaming radiation. These cases are widely studied in the literature for the physics relevant for the CMB and large scale structures. The conventional definition of isocurvature is based on setting $\zeta=0$ and a certain $\mathcal{S}_{\alpha r}\neq 0$ [Bucher:1999re]. As argued in section 2, we choose $\mathcal{S}_{\alpha\gamma}\neq 0$ instead of $\mathcal{S}_{\alpha r}\neq 0$ to fix isocurvature initial conditions. With this definition, CDI and BDI are the same, but NDI and DRDI are slightly different than those in the literature. We also list the adiabatic initial conditions for comparison.
@@
$\sigma_{\nu}$ $=$ $\frac{1}{6(15+2R_{\nu})}\omega_{b}k^{2}\tau^{3}~,$ (B.2)
@@
$\sigma_{\nu}$ $=$ $\frac{1}{6(15+2R_{\nu})}\omega_{c}k^{2}\tau^{3}$ (B.3)
@@
$\delta_{\gamma}$ $=$ $\delta_{\nu}=\frac{4}{3}\delta_{b}=\frac{4}{3}\delta_{c}=-\frac{R_{\rm dr}(11+8R_{\rm dr}+8R_{\nu})}{15+4R_{\rm dr}+R_{\nu}}$
@@
$\tilde{\sigma}_{\nu}$ $=$ $\frac{2}{3(15+4R_{\nu})}k^{2}\tau^{2}~.$ (B.10)
@@
$\tilde{\sigma}_{\nu}$ $=$ $\frac{1}{6(15+2R_{\nu})}\omega_{b}k^{2}\tau^{3}~,$ (B.11)
@@
$\tilde{\sigma}_{\nu}$ $=$ $\frac{1}{6(15+2R_{\nu})}\omega_{c}k^{2}\tau^{3}~,$ (B.12)

### v2

@@ Acknowledgement
We thank Lang Liu and Soubhik Kumar for useful discussions on this work and feedback on a draft of this manuscript. This work was supported in part by the National Key R&D Program of China (2021YFC2203100), by the National Natural Science Foundation of China (12433002, 12261131497), by CAS young interdisciplinary innovation team (JCTD-2022-20), by 111 Project (B23042), by CSC Innovation Talent Funds, by USTC Fellowship for International Cooperation, and by USTC Research Funds of the Double First-Class Initiative. PD is supported by the National Natural Science Foundation of China (Grants No. T2388102 and 12575111). JZ is supported by the National Natural Science Foundation of China (125B1023).
@@ Appendix A Evolution Functions for Primordial Perturbations
In this appendix, we show the set of Einstein equations and Boltzmann equations that govern the evolution of primordial perturbations. In the Newtonian gauge with the notation defined in section 2, the Einstein equations can be written as [54, 55, 56]
@@
To obtain the Boltzmann equations for density perturbations, we consider baryons are strongly coupled with the SM radiation bath (denoted as $\gamma_{*}$) and use the tight-coupling approximation: setting $\theta_{b}=\theta_{\gamma_{*}}$ and neglecting the anisotropic stress $\sigma_{\gamma_{*}}=0$ [56, 2]. With this approximation, the evolution functions of $\theta_{b}$ and $\theta_{\gamma_{*}}$ can be combined into one equation. In the radiation dominated era ($w=1/3$), the Boltzmann equations for the density perturbations of different species read
@@
$\theta_{c}^{\prime}$ $=$ $-\mathcal{H}\theta_{\rm c}+k^{2}\Phi~,$ (A.9)
@@
For DRDI with free-streaming dark radiation, we need to add an additional DR species which contains net anisotropic stress. The complete Boltzmann hierarchy for DR involves an arbitrarily large $\ell$ moment (typically denoted as $F_{\ell}$, see Eq. (C.7)):
@@
$F_{\rm dr\,\ell}^{\prime}$ $=$ $\frac{k}{2\ell+1}\left[\ell F_{\rm dr\,\ell-1}-(\ell+1)F_{\rm dr\,\ell+1}\right]~~~~(\ell\geq 2),$ (A.12)
where $\delta=F_{0}$, $\theta=(3/4)kF_{1}$, $\sigma=(1/2)F_{2}$. The effects of high $\ell$ moments on observables are suppressed by powers of $k\tau$ in the superhorizon limit. The Boltzmann hierarchy can therefore be simplified by truncating it at a certain $\ell_{\rm max}$ depending on the desired precision. In this study, we choose $\ell_{\rm max}=8$ and obtain closed differential equations using $F_{\ell_{\rm max}+1}\approx(2\ell_{\rm max}+1)F_{\ell_{\rm max}}/(k\tau)-F_{\ell_{\rm max}-1}$ [56] and initial conditions $F_{\ell\geq 3}=0$.
@@ Appendix B Isocurvature Initial Conditions After Neutrino Decoupling
Here we show isocurvature initial conditions after neutrino decoupling, where neutrinos are treated as free-streaming radiation. These cases are widely studied in the literature for the physics relevant for the CMB and large scale structures. The conventional definition of isocurvature is based on setting $\zeta=0$ and a certain $\mathcal{S}_{\alpha r}\neq 0$ [2]. As argued in section 2, we choose $\mathcal{S}_{\alpha\gamma}\neq 0$ instead of $\mathcal{S}_{\alpha r}\neq 0$ to fix isocurvature initial conditions. With this definition, CDI and BDI are the same, but NDI and DRDI are slightly different than those in the literature. We also list the adiabatic initial conditions for comparison.
@@
$\sigma_{\nu}$ $=$ $-\frac{1}{6(15+2R_{\nu})}\omega_{b}k^{2}\tau^{3}~,$ (B.2)
@@
$\sigma_{\nu}$ $=$ $-\frac{1}{6(15+2R_{\nu})}\omega_{c}k^{2}\tau^{3}~,$ (B.3)
@@
$\delta_{\gamma}$ $=$ $\delta_{\nu}=\frac{4}{3}\delta_{b}=\frac{4}{3}\delta_{c}=-\frac{R_{\rm dr}(11+8R_{\rm dr}+8R_{\nu})}{15+4R_{\rm dr}+4R_{\nu}}$
@@
$\tilde{\sigma}_{\nu}$ $=$ $\frac{-2}{3(15+4R_{\nu})}k^{2}\tau^{2}~.$ (B.10)
@@
$\tilde{\sigma}_{\nu}$ $=$ $-\frac{1}{6(15+2R_{\nu})}\omega_{b}k^{2}\tau^{3}~,$ (B.11)
@@
$\tilde{\sigma}_{\nu}$ $=$ $-\frac{1}{6(15+2R_{\nu})}\omega_{c}k^{2}\tau^{3}~,$ (B.12)
@@
Appendix C Second-order Perturbations of Free-streaming Dark Radiation
The energy momentum tensor for a certain species can be determined by its phase space distribution $f(\tau,q,\mathbf{x},\mathbf{n})$, where $\tau$ is the conformal time, $\mathbf{x}$ is the spatial coordinate, $\mathbf{q}=q\mathbf{n}$ is the comoving momentum with $q$ being the magnitude and $\mathbf{n}$ being the unit vector ($\delta_{ij}n^{i}n^{j}=1$). The time evolution of $f$ is controlled by the Boltzmann equation
$\frac{d}{d\tau}f=C[f],$ (C.1)
where $d/d\tau$ is the total derivative with respect to $\tau$ and $C[f]$ denotes the collision term. For free-streaming dark radiation (DR), the collision term vanishes. Therefore, the Boltzmann equation for DR is written as
$\frac{\partial f_{\rm dr}}{\partial\tau}+\frac{\partial f_{\rm dr}}{\partial x^{i}}\cdot\frac{dx^{i}}{d\tau}+\frac{\partial f_{\rm dr}}{\partial q}\cdot\frac{dq}{d\tau}+\frac{\partial f_{\rm dr}}{\partial n^{i}}\cdot\frac{dn^{i}}{d\tau}=0~.$ (C.2)
The phase space distribution of DR $f_{\rm dr}$ (up to the second order) can be decomposed as
$f_{\rm dr}(\tau,q,\mathbf{x},\mathbf{n})=\bar{f}_{\rm dr}(q)+f^{(1)}_{\rm dr}(\tau,q,\mathbf{x},\mathbf{n})+f^{(2)}_{\rm dr}(\tau,q,\mathbf{x},\mathbf{n})~.$ (C.3)
We consider the unperturbed $\bar{f}_{\rm dr}$ is the equilibrium distribution (e.g., Fermi-Dirac distribution) which is only a function of $q$. The superscript $(n)$ denotes the perturbation of the $n$-th order. Similarly, metric perturbations in Eq. (2.1) can be written as $\Phi=\Phi^{(1)}+\Phi^{(2)}$$\Psi=\Psi^{(1)}+\Psi^{(2)}$, $h_{ij}=h_{ij}^{(2)}$. Here we consider the tensor fluctuation is sourced by the product of first-order scalar perturbations, which appears at the second order.
The Boltzmann equation governs the evolution of the first-order perturbation reads
$\frac{\partial f_{\rm dr}^{(1)}}{\partial\tau}+\frac{\partial f_{\rm dr}^{(1)}}{\partial x^{i}}n^{i}=-q\frac{\partial\bar{f}_{\rm dr}}{\partial q}\left(-\frac{\partial\Phi^{(1)}}{\partial x^{j}}n^{j}+\Psi^{(1)\prime}\right),$ (C.4)
where the prime denotes the partial derivative with respect to $\tau$. The general solution to $f_{\rm dr}^{(1)}$ in the Fourier space can be obtained as
$f_{\rm dr}^{(1)}(\tau,q,\mathbf{k},\mathbf{n})$ $=$ $C(\mathbf{k},q)e^{-i\mathbf{k}\cdot\mathbf{n}\,\tau}$
[…]
where the first line comes from the solution without metric perturbations and $C(\mathbf{k},q)\approx\bar{f}_{\rm dr}(q)c(\mathbf{k})$ for $R_{\rm dr}\ll 1$ is determined by the definition of density perturbations in Eqs. (2.9) and (2). For DRDI with $R_{\rm dr}\ll 1$, the first line dominates because $\Phi^{(1)},\Psi^{(1)}$ vanishes as $R_{\rm dr}\to 0$. Therefore, we will only keep the first line for later analysis.
The Boltzmann equation for the second-order perturbation $f_{\rm dr}^{(2)}$ is [79, 78]
$\frac{\partial f_{\rm dr}^{(2)}}{\partial\tau}+\frac{\partial f_{\rm dr}^{(2)}}{\partial x^{i}}n^{i}=$ $-\frac{\partial f_{\rm dr}^{(1)}}{\partial x^{i}}n^{i}\left(\Phi^{(1)}+\Psi^{(1)}\right)-q\frac{\partial f_{\rm dr}^{(1)}}{\partial q}\left(-n^{j}\frac{\partial\Phi^{(1)}}{\partial x^{j}}+\Psi^{(1)\prime}\right)$
$-\frac{\partial f_{\rm dr}^{(1)}}{\partial n^{i}}\left[n^{i}n^{j}\frac{\partial}{\partial x^{j}}\left(\Phi^{(1)}+\Psi^{(1)}\right)-\frac{\partial}{\partial x_{i}}\left(\Psi^{(1)}+\Phi^{(1)}\right)\right]+$... (C.6)
where we have neglected terms that vanish after applying the polarization tensor $e^{ij}_{\mathbf{k},\lambda}$ because they do not source GWs. We have also dropped the term involving $h^{(2)}_{ij}$ because it leads to a damping term which is further suppressed by $R_{\rm dr}$ [80, 79, 81].
It is useful to define the brightness function of DR
$F^{(n)}_{\rm dr}=\frac{\int q^{3}\mathrm{d}qf^{(n)}_{\rm dr}}{\int q^{3}\mathrm{d}q\bar{f}_{\rm dr}}~.$ (C.7)
We note that the definition of $F_{\rm dr}^{(1)}$ matches the usual definition of $F_{\rm dr}$ in Ref. [56]. Since the dependence of $\mathbf{n}$ in first-order perturbations is only through $\mathbf{\hat{k}}\cdot\mathbf{n}$ [56], we can expand it using Legrendre polynomials $P_{\ell}$ as
$F^{(1)}_{\rm dr}(\tau,\mathbf{k},\mathbf{\hat{k}}\cdot\mathbf{n})=\sum_{\ell}^{\infty}(-i)^{\ell}(2\ell+1)F_{\textrm{dr}\,\ell}^{(1)}(\tau,\mathbf{k})P_{\ell}(\hat{\mathbf{k}}\cdot\mathbf{n}).$ (C.8)
Plugging Eq. (C) into Eq. (C.7), we obtain
$F^{(1)}_{\rm dr}(\tau,\mathbf{k},\mathbf{\hat{k}}\cdot\mathbf{n})=c(\mathbf{k})e^{-i\mathbf{k}\cdot\mathbf{n}\,\tau}~;~F^{(1)}_{\textrm{dr}\,\ell}(\tau,\mathbf{k})=c(\mathbf{k})j_{\ell}(k\tau)~~~~(\textrm{DRDI~with}~R_{\rm dr}\ll 1),$ (C.9)
where $c(\mathbf{k})$ is defined in Eq. (2.9).
To get the second-order perturbation $F^{(2)}_{\rm dr}$, we replace $f^{(2)}_{\rm dr}$ with $F^{(2)}_{\rm dr}$ in Eq. (C.6) and take the Fourier transform:
$F_{\rm dr}^{(2)\prime}(\tau,\mathbf{k},\mathbf{n})+i\mathbf{k}\cdot\mathbf{n}F^{(2)}_{\rm dr}(\tau,\mathbf{k},\mathbf{n})=S_{\rm dr}^{(2)}(\tau,\mathbf{k},\mathbf{n})~,$ (C.10)
where
$S^{(2)}_{\rm dr}(\tau,\mathbf{k},\mathbf{n})$ $=\int\frac{\mathrm{d}^{3}\mathbf{q}}{(2\pi)^{3}}\left[-(i\mathbf{q}\cdot\mathbf{n})F^{(1)}_{\textrm{dr}\,\mathbf{q}}(\Phi^{(1)}_{\mathbf{k}_{2}}+\Psi^{(1)}_{\mathbf{k}_{2}})+4F^{(1)}_{\textrm{dr}\,\mathbf{q}}\left((-i\mathbf{k}_{2}\cdot\mathbf{n})\Phi^{(1)}_{\mathbf{k}_{2}}+\Psi^{(1)\prime}_{\mathbf{k}_{2}}\right)\right$.
$\quad\quad\quad\quad\left.-\frac{\partial F^{(1)}_{\textrm{dr}\,\mathbf{q}}}{\partial n^{j}}\left(n^{j}(i\mathbf{k}_{2}\cdot\mathbf{n})(\Phi^{(1)}_{\mathbf{k}_{2}}+\Psi^{(1)}_{\mathbf{k}_{2}})-ik_{2}^{j}(\Phi^{(1)}_{\mathbf{k}_{2}}+\Psi^{(1)}_{\mathbf{k}_{2}})\right)\right$]~, (C.11)
where $\mathbf{k}_{2}=\mathbf{k}-\mathbf{q}$. Since the angular dependence of $F^{(2)}_{\rm dr}$ is controlled by two degrees of freedom $\mathbf{n}$, we shall expand it by spherical harmonics $Y_{\ell,m}$ instead of $P_{\ell}$:
$F_{\rm dr}^{(2)}(\tau,\mathbf{k},\mathbf{n})=\sum_{\ell}\sum^{\ell}_{m=-\ell}F_{\textrm{dr}\,\ell,m}^{(2)}(\tau,\mathbf{k})(-i)^{\ell}\sqrt{\frac{4\pi}{2\ell+1}}Y_{\ell,m}(\mathbf{n}).$ (C.12)
As we shall see later, the source term of induced gravitational waves depends on $F_{\textrm{dr}\,2,\sigma}^{(2)}$ with $\sigma=\pm 2$ denoting two polarizations. According to Eq. (C.11), the solution of $F_{\textrm{dr}\,2,\sigma}^{(2)}$ is
$F^{(2)}_{\textrm{dr}\,2,\sigma}(\tau,\mathbf{k})=\int_{0}^{\tau}\mathrm{d}\tilde{\tau}\ S^{(2)}_{\textrm{dr}\,2,\sigma}(\tilde{\tau},\mathbf{k})\left[j_{0}(k(\tilde{\tau}-\tau))+\frac{10}{7}j_{2}(k(\tilde{\tau}-\tau))+\frac{3}{7}j_{4}(k(\tilde{\tau}-\tau))\right],$ (C.13)
where
$S_{\textrm{dr}\,2,\sigma}^{(2)}(\tau,\mathbf{k})=-\sqrt{\frac{5}{4\pi}}\int d\Omega_{n}\ S^{(2)}_{\rm dr}(\tau,\mathbf{k},\mathbf{n})Y^{*}_{2,\sigma}(\mathbf{n}).$ (C.14)
With useful math identities and angular integrals in section C.2, we can obtain
$S^{(2)}_{{\rm dr}\,2,\sigma}=\int\frac{d^{3}\mathbf{q}}{(2\pi)^{3}}q_{i}q_{j}\sqrt{\frac{2}{3}}e^{ij}_{\mathbf{k},-\sigma}(16v_{{\rm dr}\,\mathbf{q}}\Phi_{\mathbf{k}_{2}})~~~~(\textrm{DRDI~with}~R_{\rm dr}\ll 1),$ (C.15)
where $kv_{\rm dr}=(3/4)F^{(1)}_{\textrm{dr}\,1}$ and $e^{ij}_{\mathbf{k},\sigma}$ is the polarization tensor that projects the $m=\sigma$ mode. We used $n_{i}n_{j}e^{ij}_{\mathbf{k},-\sigma}=\sqrt{\frac{2}{3}}\sqrt{\frac{4\pi}{5}}Y_{2,\sigma}^{*}(\mathbf{n})$, where we choose the $z$ direction parallel to $\mathbf{k}$ and two polarization vectors as
[…]
$e^{ij}_{\mathbf{k},2}=e^{i}(\mathbf{k})e^{j}(\mathbf{k})~,~e^{ij}_{\mathbf{k},-2}=\bar{e}^{i}(\mathbf{k})\bar{e}^{j}(\mathbf{k})$ (C.17)
We note that we use the polarization tensor $e^{ij}_{\mathbf{k},\sigma}$ here for convenience in the calculation with spherical harmonics. However, in the main text we define GWs with $\lambda=+,\times$ modes (see Eq. (3.6))
$e^{ij}_{\mathbf{k},+}=\frac{1}{\sqrt{2}}\left(e^{i}(\mathbf{k})e^{j}(\mathbf{k})-\bar{e}^{i}(\mathbf{k})\bar{e}^{j}(\mathbf{k})\right)$
$e^{ij}_{\mathbf{k},\times}=\frac{1}{\sqrt{2}}\left(e^{i}(\mathbf{k})\bar{e}^{j}(\mathbf{k})+\bar{e}^{i}(\mathbf{k})e^{j}(\mathbf{k})\right)$ (C.18)
C.1 Second-order Energy Momentum Tensor
The energy-momentum tensor of DR can be written in terms of the phase space distribution function $f_{\rm dr}$
$T^{\;\mu}_{\textrm{dr}\,\nu}=g_{\rm dr}\int\frac{\mathrm{d}P^{3}}{(2\pi)^{3}}(-g)^{-1/2}\frac{P_{\mu}P_{\nu}}{P^{0}}f_{\rm dr},$ (C.19)
where $g_{\rm dr}$ denotes the degree of freedom of DR. The second-order perturbation $\delta T^{(2)}_{\rm dr}{}_{j}^{i}$ is then given as [79, 81]
$\delta T^{(2)}_{\rm dr}{}_{j}^{i}=\frac{a^{-4}g_{\rm dr}}{(2\pi)^{3}}\int d\Omega_{n}\ q^{2}dq\ q\ n^{i}n_{j}f^{(2)}_{\rm dr}=\frac{\bar{\rho}_{\rm dr}}{4\pi}\int d\Omega_{n}n^{i}n_{j}F^{(2)}_{\rm dr}$, (C.20)
where $\bar{\rho}_{\rm dr}\equiv a^{-4}g_{\rm dr}/(2\pi)^{3}\int d^{3}\mathbf{q}\,q\bar{f}_{\rm dr}(q)$ and $F^{(2)}_{\rm dr}$ is defined in Eq. (C.7). Using the definition in Eq. (3.2) of the main text, we obtain
$\Pi_{\textrm{dr}\,ij}^{(2)}=\int\frac{d\Omega_{n}}{4\pi}n_{i}n_{j}F_{\rm dr}^{(2)}.$ (C.21)
The source term for GWs from DR component involves a term that is proportional to $\Pi_{\textrm{dr}\,ij}^{(2)}e^{ij}_{\mathbf{k},-\sigma}$, which is given as
$\Pi_{\textrm{dr}\,ij}^{(2)}e^{ij}_{\mathbf{k},-\sigma}=-\frac{1}{5}\sqrt{\frac{2}{3}}F^{(2)}_{{\rm dr}\,2,\sigma}~$, (C.22)
where we have used the relation $n_{i}n_{j}e^{ij}_{\mathbf{k},\mp 2}=\sqrt{\frac{2}{3}}\sqrt{\frac{4\pi}{5}}Y_{2,\pm 2}^{*}(\mathbf{n})$ and
$F^{(2)}_{{\rm dr}\,2,\sigma}(\tau,\mathbf{k})=\int d\Omega_{n}(-1)\sqrt{\frac{5}{4\pi}}Y_{2,\sigma}^{*}(\mathbf{n})F^{(2)}_{\rm dr}(\tau,\mathbf{k},\mathbf{n})~.$ (C.23)
Comparing Eqs. (C.13) and (C.22), we can obtain
${\Pi}^{(2)}_{\textrm{dr}\,ij}(\mathbf{k},\tau)$ $=-\int\frac{\mathrm{d}^{3}\mathbf{q}}{(2\pi)^{3}}{q}_{i}{q}_{j}\int_{0}^{\tau}\mathrm{d}\tilde{\tau}\frac{32}{15}v_{\rm dr}(\mathbf{q},\tilde{\tau})\Phi(\mathbf{k}-\mathbf{q},\tilde{\tau})~~~~~~~~~~(\textrm{DRDI with}~R_{\rm dr}\ll 1)$
$\times\left[j_{0}(k(\tilde{\tau}-\tau))+\frac{10}{7}j_{2}(k(\tilde{\tau}-\tau))+\frac{3}{7}j_{4}(k(\tilde{\tau}-\tau))\right]~.$ (C.24)
C.2 Useful Math Equations
In this section, we list a few useful math equations that relevant for calculation in this appendix.
Relations between Legrendre Polynomials with spherical harmonics:
[…]
Orthonormal conditions for spherical harmonics:
$\int\mathrm{d}\Omega_{n}Y_{\ell_{1},m_{1}}^{*}(\mathbf{n})Y_{\ell_{2},m_{2}}(\mathbf{n})=\delta_{\ell_{1}\ell_{2}}\delta_{m_{1}m_{2}}~.$ (C.26)
Integrals of the product of three spherical harmonics (Gaunt coefficients):
$\int d\Omega_{n}\,Y_{\ell_{1}m_{1}}({\mathbf{n}})Y_{\ell_{2}m_{2}}({\mathbf{n}})Y_{\ell_{3}m_{3}}({\mathbf{n}})$
$=$ $\sqrt{\frac{(2\ell_{1}+1)(2\ell_{2}+1)(2\ell_{3}+1)}{4\pi}}\times\left(\begin{array}[]{ccc}\ell_{1}&\ell_{2}&\ell_{3}\\ 0&0&0\end{array}\right)\left(\begin{array}[]{ccc}\ell_{1}&\ell_{2}&\ell_{3}\\ m_{1}&m_{2}&m_{3}\end{array}\right),$ (C.31)
where ()(...) denotes the $3j$ symbol.